(Use an Excel spreadsheet) create a household budget showing 2 columns: Monthly Budget and an Annualized Budget. Create a fictitious profile and be sure it has appropriate formatting utilized within excel, utilization of formulas when appropriate, does budget category/line items make sense, and were all items that tend to impact a budget for an average household accounted for.

***Please show the FORMULAS and where to put them on EXCEL and steps.***

*****your model should contain formulas in relevant places on EXCEL****.
Answer #1
Personal and Family Expenses
Category Monthly Budget
Current Yearly
Alimony $      576 $     6,914.75
Bank Charges $        10 $        120.00
Books/Magazines $        20 $        240.00
Business Expense $   1,152 $   13,829.50
Care of Parent/Other $      934 $   11,207.60
Cash — Miscellaneous $   1,200 $   14,400.00
Cell Phone $        15 $        180.00
Charitable Donations $      100 $     1,200.00
Child Activities $      150 $     1,800.00
Child Allowance/Expense $      200 $     2,400.00
Child Care $      150 $     1,800.00
Child Support $      100 $     1,200.00
Child school help $      450 $     5,400.00
Clothing — Client $        50 $        600.00
Clothing — Co-Client $        55 $        660.00
Clothing — Children $        50 $        600.00
Club Dues $      100 $     1,200.00
Credit Card Debt $   1,500 $   18,000.00
Dining $      600 $     7,200.00
Entertainment $      200 $     2,400.00
Gifts $        50 $        600.00
Groceries $      250 $     3,000.00
Health Care - Dental $        10 $        120.00
Health Care - Medical $        50 $        600.00
Health Care - Prescription $        50 $        600.00
Hobbies $      100 $     1,200.00
Household Items $      600 $     7,200.00
Laundry/Dry Cleaning $      100 $     1,200.00
Personal Care $      100 $     1,200.00
Personal Loan Payment $      150 $     1,800.00
Pet Care $         -   $                -  
Public Transportation $        25 $        300.00
Recreation $      200 $     2,400.00
Self Improvement $        50 $        600.00
Student Loan $      600 $     7,200.00
Vacation/Travel $   1,000 $   12,000.00
Other: $      577 $     6,924.00
$ 11,525 $ 138,295.00
Personal and Family Expenses
Category Monthly Budget
Current Yearly
Alimony =+B41*0.05 =+B4*12
Bank Charges 10 =+B5*12
Books/Magazines 20 =+B6*12
Business Expense =+B41*0.1 =+B7*12
Care of Parent/Other =+B41*0.08+12 =+B8*12
Cash — Miscellaneous 1200 =+B9*12
Cell Phone 15 =+B10*12
Charitable Donations 100 =+B11*12
Child Activities 150 =+B12*12
Child Allowance/Expense 200 =+B13*12
Child Care 150 =+B14*12
Child Support 100 =+B15*12
Child school help 450 =+B16*12
Clothing — Client 50 =+B17*12
Clothing — Co-Client 55 =+B18*12
Clothing — Children 50 =+B19*12
Club Dues 100 =+B20*12
Credit Card Debt 1500 =+B21*12
Dining 600 =+B22*12
Entertainment 200 =+B23*12
Gifts 50 =+B24*12
Groceries 250 =+B25*12
Health Care - Dental 10 =+B26*12
Health Care - Medical 50 =+B27*12
Health Care - Prescription 50 =+B28*12
Hobbies 100 =+B29*12
Household Items 600 =+B30*12
Laundry/Dry Cleaning 100 =+B31*12
Personal Care 100 =+B32*12
Personal Loan Payment 150 =+B33*12
Pet Care 0 =+B34*12
Public Transportation 25 =+B35*12
Recreation 200 =+B36*12
Self Improvement 50 =+B37*12
Student Loan 600 =+B38*12
Vacation/Travel 1000 =+B39*12
Other: 577 =+B40*12
=(76035+62260)/12 =+B41*12